OpenStudy (anonymous):

all of the letters of the word DESPAIR are arranged in a row. If the Letters were arranged randomly, calculate the probability that: the vowels are together That the first letter is A, given that the vowels are together?

OpenStudy (shubhamsrg):

1)vowels in DESPAIR = e,a,i =3 let eai be one 1 letter..so DESPAIR has 5 letters to arrange them,,we have 5! ways but eai can also exchange places 3! times so final ans is 5! * 3! 2) now if A is first letter and the vowels remain together always,, same way,,we have 5! ways and eai can exchange places 2! times as A is fixed so final ans here is 5! * 2! hope that could help..

OpenStudy (shubhamsrg):

sorry sorry,,correction in 2nd part..

OpenStudy (shubhamsrg):

the 5 set of letters in despair can exchange places 4! times as set of vowels is fixed..and multiplying that by 2!,,ans is 4! * 2!

OpenStudy (campbell_st):

the number of arrangements of the letter is 7! Number of arrangements with vowels together is 5! x 3! P(vowels together) (5!x3!)/7! = 1/7 Vowels together 1st letter A = 5! X 2! P(A 1st letter) = (5!x2!)/7! = 1/21

OpenStudy (shubhamsrg):

and yes,,for probability,,divide each ans by 7! ..sorry i missed that earlier,,damit!

OpenStudy (shubhamsrg):

and maybe i have even misunderstood the second part..gosh,,whats happening!

OpenStudy (campbell_st):

ages since I did this stuff... always rusty with it... so get someone to check mine

OpenStudy (kinggeorge):

@campbell_st it looks all good to me except in part two you should have \(4!\cdot2!\) since you're only permuting the 4 letters d, s, p, r. and assuming that all the vowels are first.

OpenStudy (dumbcow):

i agree but it depends on the wording....does it mean A is first letter of word or just first of the vowels

OpenStudy (campbell_st):

thanks... 2nd part should be 4! x 2! P(A 1st then vowels) = 4!x2!/7! 1/105

OpenStudy (kinggeorge):

a is the first letter of the word, but the vowels are all together, so that means all the vowels must be before the rest of the letters.

OpenStudy (anonymous):

hmmmm

OpenStudy (campbell_st):

I always disliked these questions...

OpenStudy (anonymous):

the answer to the second part of this is 2!4!/ 3! 5! so 1/15 except im confused.... I understand part a though D:

OpenStudy (anonymous):

why is it 2! 4! ?

OpenStudy (anonymous):

why isnt it 2! X 5! ?

OpenStudy (campbell_st):

assume the vowels come 1st A is fixed... then there are 2 ways of arranging the remaining 2 vowels.... next comes the 4 letters D,S,P,R, the can be arranged in 4! ways

OpenStudy (campbell_st):

hence 2! x 4!

OpenStudy (anonymous):

ohh so you dont make it 5!, (5 letters) because 2! cannot be moved around anyway?